OpenStudy (anonymous):

Rachel works at a bookstore. On Tuesday, she sold twice as many books as she did on Monday. On Wednesday, she sold 6 fewer books than she did on Tuesday. During the three days Rachel sold 19 books. Create an equation that can be used to find m, the number of books Rachel sold on Monday.

OpenStudy (anonymous):

2x-6y=19 but i'm not sure

OpenStudy (anonymous):

Let M be the number of books Rachel sold on Monday, T the number of books sold on Tuesday, and W be the number of books sold on Wednesday. From your second sentence, we get T=2M, because she sold twice as many books on Tuesday than Monday. "On Wednesday, she sold 6 fewer books than she did on Tuesday." gives us W = T - 6 But we know that T = 2M, so W = 2M -6 Finally, on all three days, she sold 19 books: M + T + W = 19. Now, combine the 3 equations! M + T + W = 19 M + 2M + 2M-6 = 19 5M = 25 M = 5